# Magnetic effects in the Hadron Resonance Gas

### v1

@@ Abstract
We discuss the modeling of the hadronic phase of QCD at finite magnetic field in the framework of hadron resonance gas (HRG). We focus on the statistical description of particle yields that include contribution from resonance decays. We demonstrate that the swift increase in the number of protons with magnetic field predicted in the HRG is due to the ill-defined description of higher-spin states. We discuss fluctuations of conserved charges and show that at present the qualitative comparison of the model predictions with the Lattice QCD data should be treated with care. We also discuss the principle of detailed balance which allows to study the magnetic field dependence of neutral resonances.
@@ I Introduction
Magnetic fields generated in non-central heavy-ion collisions (HICs) are among the strongest in the universe [1, 2, 3]. Since quarks and antiquarks are electrically charged and influenced by magnetic field, understanding the interplay between the magnetic field and the strong interactions governed by quantum chromodynamics (QCD) is crucial for properly comprehending non-central heavy-ion collisions. Understanding the impact of an extremely strong magnetic field on the strongly interacting matter is also relevant for studying magnetars [4] and the early universe [5]. The magnetic field induces a variety of interesting phenomena in the QCD matter, such as anomaly-induced transport phenomena e.g., Chiral Magnetic Effect (CME) [6] and Chiral Vortical Effect (CVE) [7], charged $\rho$ meson condensation [8], modification of QCD Debye mass [9], QCD equation of state (EoS) [10, 9, 11, 12], dilepton production rate [13, 14].
@@
Investigation of QCD in the background of magnetic fields is also interesting from the theoretical standpoint since it can be simulated from the first principle lattice QCD (LQCD) methods even for large fields in contrast to finite baryon chemical potential where the simulations are marred by the numerical sign problem. LQCD simulations indicate a nontrivial interplay between the magnetic field and strong interactions, such as magnetic catalysis at low temperatures [15, 16], and inverse magnetic catalysis [17, 18, 19] which influences the QCD phase diagram. These effects have also been studied widely using several other formalisms viz., Nambu-Jona-Lasinio (NJL) model, Polyakov improved NJL (PNJL) model [20], quark-meson (QM) model [20], gauge-gravity duality [21]. However, the latter effect is not captured in commonly used chiral models [22]. It can, nevertheless, be generated by considering additional in-medium effects (see e.g. [23]). It has also been found that the pseudocritical temperature of the QCD crossover decreases with the magnetic field (see, e.g., Refs. [24, 25, 18]), with the possible existence of the critical point at large field intensity [25, 26]. For recent reviews of QCD in external magnetic field, see [22, 27, 28, 29]. Recently, there has been significant attention to the fluctuations of conserved charges in the magnetic field background using LQCD methods [30, 31, 12, 32]. It has been observed that a finite magnetic field has a nontrivial effect on these observables. They are of direct phenomenological interest because they are sensitive to in-medium degrees of freedom and can be measured experimentally.
@@
The hadron resonance gas (HRG) model has been successful in describing the LQCD data on the EoS and fluctuations of conserved charges at vanishing magnetic field [33, 34] as well as the hadron yields from heavy-ion collisions [35]. In [36], the HRG model has been generalized to include the effects of finite magnetic field. The model has been used to study the impact of magnetic field on the QCD EoS [36, 10], fluctuations of conserved charges [37, 38, 31, 39] as well as other effects, such as transport coefficients [40, 41]. The behavior of spin-00 and spin-$1/2$ particles in an external magnetic field is well understood and involves the Landau quantization in the directions perpendicular to the magnetic field [42]. However, when the same prescription is applied to higher-spin particles, it leads to imaginary dispersion relations at strong fields as well as instabilities in the pressure [36]. This suggests that the treatment of higher-spin particles in the HRG model is doubtful and requires further investigation.
@@
In this work, we study the fluctuations of conserved charges and particle yields at a finite magnetic field using the HRG model. We find that the magnetic field enhances the baryon yields, while $\pi^{\pm}$ and $K^{\pm}$ yields are weakly affected. With our choice of resonances included in HRG, we find that the model underestimates the LQCD baryon number fluctuations at a weak magnetic field while overestimating them at large $eB$, the latter originating from the divergences induced by higher-spin resonances. We also gauge the impact of the magnetic field on neutral resonances by implementing certain detailed balance conditions. We compare the proton and pion yields obtained using HRG and our detailed balance requirement and observe a substantial difference, indicating the need for further research on the impact of strong magnetic field on resonance properties.
@@ II Hadronic Matter in finite magnetic field
$f_{i}=\frac{1}{e^{\beta(\epsilon_{i}-\mu_{i})}\mp 1}$ (3)
@@
where $s_{z}$ is the $z$-component of the particle’s spin and $l\in\{0,1,2,\ldots\}$ numbers the Landau levels. We note that the dispersion relation in Eq. (4) is exact for structureless spin-00 and spin-$1/2$ particles [22]. It has also been found that spin-$3/2$ particles, governed by the Rarita-Schwinger equation, exhibit a non-causal behavior in the presence of external magnetic field [45, 46, 47], which can be fixed by adding non-minimal interaction term to the Lagrangian [48, 49] and end up with a similar form of dispersion relation depicted at Eq. (4).
We stress that Eq. (4) neglects the compositeness of the state. Theoretical issues can arise when the scale of magnetic field resolves the structure of hadron states, i.e., $\sqrt{|Q|B}>m_{\pi}$. The problem becomes more apparent for higher spin states, where the mass of particles can be substantially reduced and eventually leads to a complex dispersion relation (signaling instability). Resolving these issues would require a more sophisticated treatment of the decay dynamics [50, 51].
The phase-space integral of a charged particle acquires the following form [52, 53]
@@
where $\gamma$ is the spin degeneracy factor and $Q$ is the electric charge of the particle. As a result, the thermal part of the pressure becomes
$P_{i}=\pm T\frac{|Q_{i}|B}{2\pi^{2}}\sum_{s_{z}}\sum_{l=0}^{\infty}\int\limits_{0}^{\infty}\mathrm{d}p_{z}\;\log\left(1\pm f_{i}\right)\,.$ (6)
@@
We note that the pressure of a finite magnetic field contributes to the bulk thermodynamic quantities, and the vacuum part of the thermodynamic pressure must be regularized [36]. In this work, we focus on the densities and fluctuations of conserved charges which do not depend on the vacuum part of the pressure. Thus, we omit it in our considerations. We also note that the magnetic field does not distinguish between positive and negative electric charges (cf. Eqs. (4) and (6) ).
@@
where $\langle N_{i}\rangle$ is an average number of particle $i$ in volume $V$. To quantify the change of the medium composition due to the presence of a finite magnetic field, we study the normalized densities of selected hadrons at fixed temperature $T=0.155~{}\rm GeV$. This is depicted in Fig. 1. The densities of $p$, $\Sigma^{+}$, and $\Xi^{-}$ generally increase with the magnetic field. This is attributed to the lowest Landau level being magnetic field-independent and hence its contribution grows linearly with $eB$ (cf. Eqs. (4) and (6)). In contrast to this, the densities of $\pi^{-}$ and $K^{-}$ are exponentially suppressed because in this case, all Landau levels increase with magnetic field. We note that there is a clear mass ordering for spin-$1/2$ baryons and spin-00 mesons. The exceptions to mass ordering are the higher-spin particles, $\rho^{+}$ and $\Omega^{-}$, whose increase is much greater than for other particle species. This swift increase is caused by the contribution of the lowest Landau level being negative for these particles, see Eq.(4). As mentioned, the extrapolation of Eq. (4) to higher-spin states requires the assumption of structureless hadrons, which can be questionable depending on the size of magnetic field and resonance structure of the state in question.
@@
where $N^{\rm th}_{i}=Vn_{i}$ is the average thermal number defined through Eq. (7), the sum in the second term is overall decays with particle $i$ in the final state, and $\Gamma_{r\rightarrow i}$ is the branching ratio of the resonance $r$ to particle $i$.
In Fig. 2, we show the particle yields normalized to the yields at the vanishing magnetic field. At small values of $eB$, the $\pi^{+}$ and $K^{+}$ yields decrease with increasing $eB$, but at a stronger magnetic field, their yields increase with increasing $eB$. Nevertheless, at $eB\lesssim 0.2~{}\rm GeV^{2}$, $\pi^{+}$ and $K^{+}$ yields change only a few percent. However, at the same values of $eB$, baryon yields increase substantially. The largest increase is seen for protons, due to decaying $\Delta$ resonances. We warn the reader again about the uncertainties related in the description of high spin ($s\geq 1$) states and note that the weak dependency of $\Sigma$ and $\Xi$ yields on the magnetic field may be a result of the way fewer known strange than non-strange resonances.
In the literature, the observed particle yields are often discussed in terms of particle ratios instead of actual yields to avoid dependency on system size. To facilitate the discussion in terms of particle ratios, we show the particle yields normalized to the yield of $\pi^{+}$ in Fig. 3. Again, the prominent feature is the substantial increase of the $p/\pi$ ratio with the increasing strength of the magnetic field. We note that even if the $\Omega^{+}$ yield increases way faster than the pion yield in Fig. 2, it is still so tiny compared to the pion yield that the change in the ratio cannot be seen at the scale of Fig. 3.
@@
We note that the coefficient $c_{2}$ is evaluated at vanishing chemical potentials. At a sufficiently small net-baryon density, the strangeness neutrality condition can be approximated by $\mu_{S}=0$. Thus, one can expect that $\chi_{BB}\approx 2c_{2}$, as is found to be the case in the HRG model. This allows us to compile the LQCD data on $\chi_{BB}$ [31] and $c_{2}$ [32] in a single figure. They are shown in Fig. 4 for $T=0.145$ and 0.1550.1550.155~{}0.155GeV. We find that the HRG model systematically underestimates the LQCD data for $eB\leq 0.16~{}\rm GeV^{2}$ at both temperatures. We note that this is in contrast to the HRG model with an augmented list of resonances which includes states predicted by relativistic quark models. Such a HRG model reproduces $\chi_{BB}$ at vanishing magnetic field [34], but, as the magnetic field increases, both the HRG model with established states only and the model with all the predicted states, start to overestimate the LQCD results. The calculated susceptibility increases rapidly and drastically overshoots the LQCD data at $eB=0.3~{}\rm GeV^{2}$. The contribution of baryons with $s\geq 3/2$ ranges from $75\%$ at vanishing $eB$ to $95\%$ at $eB=0.3~{}\rm GeV^{2}$. This result exposes the weakness of the current treatment of the higher-spin resonances.
Ultimately, to fully identify the contributions of higher-spin states to thermodynamic observables their internal structure should be taken into account. This, however, requires a dynamical treatment of resonances, e.g. the S-matrix formulation of statistical mechanics [58, 59, 60], suitably extended to the presence of $eB$, and is currently under development. As a temporary attempt, we consider the detailed balance approach to neutral particles (see the next Section).
Our major observation is that the current uncertain treatment of high-spin particles in a finite magnetic field prevents direct comparison of the HRG model with the experimental data from HIC and results obtained in LQCD simulations.
@@ III Detailed Balance
As mentioned, we assume neutral particles to have zero magnetic moments. Consequently, they are not affected by the magnetic field at all. However, in chemically equilibrated HRG detailed balance prevails, e.g., the decay rate of rho mesons is equal to the scattering rate of pions forming $\rho$ resonances. The scattering rate depends on the densities and, as seen in Fig. 1, pion density decreases with the magnetic field. Thus, the $\rho^{0}$ density may depend on the magnetic field after all.
It is known that a weak magnetic field causes a negligibly small change in the width of $\rho^{0}$ [50]. Thus in the presence of a magnetic field, the $\rho^{0}$ decay rate changes only if the $\rho^{0}$ density changes, and the change in rate is proportional to the change in density.
Unfortunately, we cannot evaluate the scattering rate of pions using the conventional kinetic theory calculation, since particles in magnetic field do not have well-defined momenta. To gain insight into how the detailed balance requirement could affect the densities of neutral resonances, we make the bold assumption that the magnetic field affects the pion scattering rate (and thus $\rho^{0}$ production rate) only by changing their densities. In other words, at fixed temperature,
@@
Detailed balance must prevail not only between neutral resonances and their daughters but also between charged resonances and their daughters. To gauge the overall uncertainty in the chemical composition of HRG in magnetic field, we assume that $\Delta$ and $\rho$ densities are not given by Eqs. 47, but all charged states of $\Delta$ and $\rho$ gain effective chemical potentials similar to $\Delta^{0}$ and $\rho^{0}$. We keep all the other densities unchanged and evaluate the pion and proton yields after decays. The resulting yields and their ratios are shown in Fig. 6. As expected, the reduction in $\rho$ and $\Delta$ densities leads to reduced yields, and since the reduction in $\rho$ is larger than in $\Delta$, the proton to pion ratio is larger.
In the literature, the centrality dependence of proton to pion ratio has been suggested as a measure of the magnetic field in heavy-ion collisions, with the assumption that the field is negligible in most central collisions, and largest in most peripheral. The ALICE collaboration has measured the $p/\pi$ ratio in Pb+Pb collisions at $\sqrt{s_{NN}}=5.02~{}\rm TeV$ [61], and reported that the ratio increases by a factor $\sim 1.3$ from central to peripheral collisions. In the lower panel of Fig. 6, we have included dots on the $p/\pi$ curves indicating where the ratio has increased by 30% compared to its value at vanishing magnetic field. We see that our assumptions about the chemical composition of HRG lead to similar $\sim 30$% uncertainty in the allowed strength of the magnetic field.

### v2

@@ Abstract
We discuss the modeling of the hadronic phase of QCD at finite magnetic field in the framework of hadron resonance gas (HRG). We focus on the statistical description of particle yields that include contribution from resonance decays. We demonstrate that the swift increase in the number of protons with magnetic field predicted in the HRG is due to the assumption of structureless resonances. We discuss fluctuations of conserved charges and show that at present the qualitative comparison of the model predictions with the Lattice QCD data should be treated with care. We also discuss the principle of detailed balance which allows us to study the magnetic field dependence of neutral resonances.
@@ I Introduction
Magnetic fields generated in non-central heavy-ion collisions (HICs) are among the strongest in the universe [1, 2, 3]. Since quarks and antiquarks are electrically charged and influenced by magnetic field, understanding the interplay between the magnetic field and the strong interactions governed by quantum chromodynamics (QCD) is crucial for properly comprehending non-central HICs. Understanding the impact of an extremely strong magnetic field on the strongly interacting matter is also relevant for studying magnetars [4] and the early universe [5]. The magnetic field induces a variety of interesting phenomena in the QCD matter, such as anomaly-induced transport phenomena, e.g., Chiral Magnetic Effect (CME) [6] and Chiral Vortical Effect (CVE) [7], charged $\rho$ meson condensation [8] and modifications of QCD Debye mass [9], QCD equation of state (EoS) [10, 9, 11, 12] and dilepton production rate [13, 14].
@@
Investigation of QCD in the background of magnetic fields is also interesting from the theoretical standpoint since it can be simulated from the first principle lattice QCD (LQCD) methods even for large fields in contrast to finite baryon chemical potential where the simulations are marred by the numerical sign problem. LQCD simulations indicate a nontrivial interplay between the magnetic field and strong interactions, such as magnetic catalysis at low temperatures [15, 16], and inverse magnetic catalysis [17, 18, 19] which influences the QCD phase diagram. These effects have also been widely studied using several other formalisms, viz. Nambu-Jona-Lasinio (NJL) model, Polyakov improved NJL (PNJL) model [20], quark-meson (QM) model [20], and gauge-gravity duality [21]. However, the inverse magnetic catalysis is not captured in commonly used chiral models [22]. Nevertheless, it can be generated by considering additional in-medium effects (see e.g. [23]). It has also been found that the pseudocritical temperature of the QCD crossover decreases with increasing magnetic field (see, e.g., Refs. [24, 25, 18]), with the possible existence of a critical point at large field strength [25, 26]. For recent reviews of QCD in external magnetic field, see [22, 27, 28, 29]. Recently, the fluctuations of conserved charges in the magnetic field background have attracted considerable interest and they have been evaluated using LQCD methods [30, 31, 12, 32]. It has been observed that a finite magnetic field has a nontrivial effect on these fluctuations. They are of direct phenomenological interest because they are sensitive to in-medium degrees of freedom and can be measured experimentally.
@@
The hadron resonance gas (HRG) model has been successful in describing the LQCD data on the EoS and fluctuations of conserved charges at vanishing magnetic field [33, 34] as well as the hadron yields from heavy-ion collisions [35]. In [36], the HRG model has been generalized to include the effects of finite magnetic field. The model has been used to study the impact of magnetic field on the QCD EoS [36, 10], fluctuations of conserved charges [37, 38, 31, 39] as well as other effects, such as transport coefficients [40, 41]. The behavior of spin-00 and spin-$1/2$ particles in an external magnetic field is well understood and involves the Landau quantization in the directions perpendicular to the magnetic field [42]. However, when the same prescription is applied to higher-spin resonances, it leads to imaginary dispersion relations at strong fields as well as instabilities in pressure [36]. This is due to the bold assumption of structureless resonance states in the HRG model. This suggests that the treatment of higher-spin resonances in the HRG model is doubtful and requires further investigation.
@@
In this work, we study the fluctuations of conserved charges and particle yields at a finite magnetic field using the HRG model. We find that the magnetic field enhances the baryon yields, while $\pi^{\pm}$ and $K^{\pm}$ yields are weakly affected. With our choice of resonances included in HRG, we find that the model underestimates the LQCD baryon number fluctuations at a weak magnetic field while overestimating them at large $eB$, the latter originating from the divergences induced by higher-spin resonances. We also gauge the impact of the magnetic field on neutral resonances by implementing detailed balance conditions. We compare the proton and pion yields obtained using HRG and our detailed balance requirement and observe a substantial difference, indicating the need for further research on the impact of strong magnetic field on resonance properties.
@@ II Hadronic Matter in finite magnetic field
$f_{i}=\frac{1}{e^{\beta(\epsilon_{i}-\mu_{i})}\pm 1}$ (3)
@@
where $s_{z}$ is the $z$-component of the particle’s spin and $l\in\{0,1,2,\ldots\}$ numbers the Landau levels. We note that the dispersion relation in Eq. (4) is exact for structureless spin-00 and spin-$1/2$ particles [22]. On the other hand, it has been found that spin-$3/2$ particles, governed by the Rarita-Schwinger equation, exhibit a non-causal behavior in the presence of external magnetic field [45, 46, 47], which can be fixed by adding a non-minimal interaction term to the Lagrangian [48, 49]. Such a procedure leads to a dispersion relation similar to Eq. (4).
We stress that Eq. (4) neglects the compositeness of resonances, that is, the possibility of decaying to daughter particles. This assumption becomes questionable when the scale of magnetic field resolves the structure of hadron states, i.e., $\sqrt{|Q|B}>m_{\pi}$. Problems become more apparent for higher spin states, where the effective mass of particles can be substantially reduced, which eventually leads to a complex dispersion relation (signaling instability). Resolving these issues would require a more sophisticated treatment of the decay dynamics [50, 51].
[…]
When the energies of particles are given by Eq. (4), the phase-space integral of a charged particle acquires the following form [64, 65]
@@
where $\gamma$ is the spin degeneracy factor and $Q$ is the electric charge of the particle. The presence of a homogeneous background magnetic field breaks the spherical symmetry, causing the pressure transverse to the magnetic field ($P^{\perp}$) to be different than the pressure parallel to it ($P^{\parallel}$). As demonstrated in Ref. [66] for a Fermi gas, the system satisfies the canonical relations $\Omega=-P^{\parallel}$, and $P^{\perp}=P^{\parallel}-BM$, where $M=-\partial\Omega/\partial B=\partial P^{\parallel}/\partial B$ is the magnetization. $P^{\parallel}$ can be obtained by applying prescription (5) to Eq. (2),
$P^{\parallel}_{i}=\pm T\frac{|Q_{i}|B}{2\pi^{2}}\sum_{s_{z}}\sum_{l=0}^{\infty}\int\limits_{0}^{\infty}\mathrm{d}p_{z}\;\log\left(1\pm f_{i}\right)\,.$ (6)
However, the contribution of the magnetization current density to the transverse pressure [67] is not considered in this formalism. If this contribution is taken into account, the pressure becomes isotropic [67] (See also Refs. [68, 69]). Nevertheless, it has been shown [70] that whether one should include this contribution or not, depends on the formal definition of pressure as a derivative of the free energy $F$, i.e., whether the magnetic field or the flux of the field is kept constant while taking the derivative. In both cases observables such as the energy density, entropy, number density, and fluctuations of conserved charges are equal and do not depend on the scheme used to compute the pressure [71]. Therefore, we use Eq. (6) to define pressure of the hadron gas and omit the $\parallel$ symbol in the remainder of this work.
@@
The pressure of a finite magnetic field contributes to the other thermodynamic quantities as well, and the vacuum part of the thermodynamic pressure must be properly renormalized [36]. In this work, we focus on the densities and fluctuations of conserved charges which do not depend on the vacuum part of the pressure. Thus, we omit it in our considerations. We also note that the magnetic field does not distinguish between positive and negative electric charges (cf. Eqs. (4) and (6)).
@@
where $\langle N_{i}\rangle$ is an average number of particle $i$ in volume $V$. To quantify the change of the medium composition due to the presence of a finite magnetic field, we study the densities (normalized to their densities at vanishing magnetic field) of selected hadrons at fixed temperature $T=0.155~{}\rm GeV$. This is depicted in Fig. 1. The densities of $p$, $\Sigma^{+}$, and $\Xi^{-}$ generally increase with the magnetic field. This is attributed to the lowest Landau level being magnetic-field independent and hence its contribution grows linearly with $eB$ (cf. Eqs. (4) and (6)). In contrast to this, the densities of $\pi^{-}$ and $K^{-}$ are exponentially suppressed because in this case the energies of all Landau levels increase with magnetic field. We note that there is a clear mass ordering for spin-$1/2$ baryons and spin-00 mesons: the lighter the particle, the larger the effect. The exceptions to this mass ordering are the higher-spin particles, $\rho^{+}$ and $\Omega^{-}$, whose increase is much greater than for other particle species. This swift increase is caused by the contribution of the lowest Landau level being negative for these particles, see Eq.(4). As mentioned, the extrapolation of Eq. (4) to higher-spin states requires the assumption of structureless hadrons, which can be questionable depending on the strength of magnetic field and the structure of the resonance in question.
@@
where $\langle N^{\rm th}_{i}\rangle=Vn_{i}$ is the average thermal number defined through Eq. (7), the sum in the second term is over all decays with particle $i$ in the final state, and $\Gamma_{r\rightarrow i}$ is the branching ratio of the resonance $r$ to particle $i$.
In Fig. 2, we show the particle yields normalized to the yields at the vanishing magnetic field. At small values of $eB$, the $\pi^{+}$ and $K^{+}$ yields decrease with increasing $eB$, but at stronger magnetic field, their yields increase with increasing $eB$. Nevertheless, at $eB\lesssim 0.2~{}\rm GeV^{2}$, $\pi^{+}$ and $K^{+}$ yields change only a few percent. However, at the same values of $eB$, proton yields increase substantially, due to decaying $\Delta$ resonances. We warn the reader again about the uncertainties related in the description of high spin ($s\geq 1$) states and note that the weak dependence of $\Sigma$ and $\Xi$ yields on the magnetic field may be a result of the way fewer known strange than non-strange resonances.
In the literature, the observed particle yields are often discussed in terms of particle ratios instead of actual yields to avoid dependency on system size. To facilitate the discussion in terms of particle ratios, we show the particle yields normalized to the yield of $\pi^{+}$ in Fig. 3. Again, the prominent feature is the behavior of protons: the $p/\pi$ ratio increases substantially with the increasing strength of the magnetic field. We note that even if the $\Omega^{+}$ yield increases way faster than the pion yield in Fig. 2, it is still so tiny compared to the pion yield that the change in the ratio cannot be seen at the scale of Fig. 3.
@@
We note that the coefficient $c_{2}$ is evaluated at vanishing chemical potentials. At a sufficiently small net-baryon density, the strangeness neutrality condition can be approximated by $\mu_{S}=0$. Thus, one can expect that $\chi_{BB}\approx 2c_{2}$, as is found to be the case in the HRG model. This allows us to compile the LQCD data on $\chi_{BB}$ [31] and $c_{2}$ [32] in a single figure. They are shown in Fig. 4 for $T=0.145$ and 0.1550.1550.155~{}0.155GeV. We find that the HRG model systematically underestimates the LQCD data for $eB\leq 0.16~{}\rm GeV^{2}$ at both temperatures. We note that this is in contrast to the HRG model with an augmented list of resonances which includes states predicted by relativistic quark models. Such a HRG model reproduces $\chi_{BB}$ at vanishing magnetic field [34], but, as the magnetic field increases, both the HRG model with established states only, and the model with all the predicted states, start to overestimate the LQCD results. The calculated susceptibility increases rapidly and drastically overshoots the LQCD data at $eB=0.3~{}\rm GeV^{2}$. The contribution of baryons with $s\geq 3/2$ ranges from $75\%$ at vanishing $eB$ to $95\%$ at $eB=0.3~{}\rm GeV^{2}$. This result exposes the weakness of the current treatment of the higher-spin resonances.
Ultimately, to fully evaluate the contributions of higher-spin states to thermodynamic observables their internal structure should be taken into account. This, however, requires a dynamical treatment of resonances, e.g. the S-matrix formulation of statistical mechanics [76, 77, 78], suitably extended to the presence of $eB$, and is currently under development. As a temporary attempt, we consider the detailed balance approach to neutral particles (see the next Section).
Based on our results and arguments presented above, we conclude that the current uncertain treatment of high-spin particles in a finite magnetic field prevents a direct comparison of the HRG model with both the experimental data from HICs and results obtained in LQCD simulations.
@@ III Detailed Balance
So far, we have assumed neutral particles to have zero magnetic moments. Consequently they were not affected by the magnetic field at all. However, in chemically equilibrated HRG detailed balance prevails, e.g., the decay rate of rho mesons is equal to the scattering rate of pions forming $\rho$ resonances. The scattering rate depends on the densities and, as seen in Fig. 1, pion density decreases with the magnetic field. Thus, the $\rho^{0}$ density may depend on the magnetic field after all.
However, systematic calculation of the decay rates of resonances is challenging. For example, the pair production in the decay process of a neutral scalar boson into two charged scalar bosons is suppressed by the presence of external magnetic field [79], a neutral boson emission rate by a fermion is enhanced by the magnetic field [80]. Nevertheless, it is known that a weak magnetic field causes a negligibly small change in the width of $\rho^{0}$ [50, 81]. Thus, we make an approximation that in the presence of a magnetic field, the $\rho^{0}$ decay rate changes only if the $\rho^{0}$ density changes, and the change in the rate is proportional to the change in density.
Unfortunately, we cannot evaluate the scattering rate of pions using a conventional kinetic theory calculation, since particles in magnetic field do not have well-defined momenta. To gain insight into how the detailed balance requirement could affect the densities of neutral resonances, we make the bold assumption that the magnetic field affects the pion scattering rate (and thus $\rho^{0}$ production rate) only by changing their densities. In other words, at fixed temperature,
@@
Detailed balance must prevail not only between neutral resonances and their daughters, but also between charged resonances and their daughters. To gauge the overall uncertainty in the chemical composition of HRG in magnetic field, we assume that the densities of $\Delta$ and $\rho$ are not given by Eqs. 47, but all charged states of $\Delta$ and $\rho$ gain effective chemical potentials similar to $\Delta^{0}$ and $\rho^{0}$. We keep all the other densities unchanged and evaluate the pion and proton yields after decays. The resulting yields and their ratios are shown in Fig. 6. As expected, the reduction in $\rho$ and $\Delta$ densities leads to reduced yields, and since the reduction in $\langle p\rangle$ is greater than in $\langle\pi^{+}\rangle$, the proton-to-pion ratio becomes smaller.
In the literature, particle ratios have been used to constrain the magnetic field in heavy-ion collisions (e.g., [82, 83]). Here we apply this idea to the proton-to-pion ratio in particular. We assume that the field is negligible in most central collisions, largest in most peripheral, and that the centrality dependence of the $p/\pi$ ratio is entirely due to the change in the magnetic field. The ALICE collaboration has measured the $p/\pi$ ratio in Pb+Pb collisions at $\sqrt{s_{NN}}=5.02~{}\rm TeV$ [84], and reported that the ratio increases by a factor $\sim 1.3$ from central to peripheral collisions. In the lower panel of Fig. 6, we have included dots on the $p/\pi$ curves indicating where the ratio has increased by 30% compared to its value at vanishing magnetic field. We find that this occurs at $eB=0.13~{}{\rm GeV^{2}}\simeq 6.6~{}m_{\pi}^{2}$ and $eB=0.20~{}{\rm GeV^{2}}\simeq 10.1~{}m_{\pi}^{2}$. Thus, our assumptions about the chemical composition of HRG lead to a $\sim 30$% uncertainty in the allowed strength of the magnetic field.